(b) Find the area of the region inside the cardioid r=2 + 2
1
Expert's answer
2021-08-16T12:33:16-0400
"r=2+2\sin \theta"

"A=\displaystyle\int_{0}^{2\pi}\dfrac{1}{2}r^2d\theta=\displaystyle\int_{0}^{2\pi}\dfrac{1}{2}(2+2\sin \theta)^2d\theta"

"=\displaystyle\int_{0}^{2\pi}(2+4\sin\theta+2\sin^2\theta)d\theta"

"=\displaystyle\int_{0}^{2\pi}(2+4\sin\theta+1-\cos(2\theta))d\theta"

"=[3\theta-4\cos\theta-\dfrac{1}{2}\sin(2\theta)]\begin{matrix}
 2\pi\\
 0
\end{matrix}"

"=6\pi-4-0-(0-4-0)=6\pi"

"Area=6\pi" square units.